OpenStudy (anonymous):

Three times a number decreased by 14 is three times the number. Find the number and justify the solution., the answer is no solution right?

OpenStudy (anonymous):

should be 3x-14=3x which leaves 14=1

OpenStudy (anonymous):

so no solution

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

3x - 14 = 3x -14 = 0 No solution

OpenStudy (anonymous):

ok so i was right thx

OpenStudy (anonymous):

Yeah, can't add or subtract a value from a number and expect the result to be the same. Makes sense logically for there to be no solution also :)